How many ways can you arrange 6 men and 9 women seated on a round table and 2 mens are not seated to each other? thank you very much
Answer by sudhanshu_kmr(1152)


There are 6 men and 9 women, restriction is with men not women
so, firstly we arrange 9 women on a round table

no. of ways = (9-1)!= 8!

now there are 8 spaces to seat men..
( a men can seat between 2 women for confirm separation like F m F - F )

no. of ways to arrange 6 men into 8 seats = 8P6


total no. of ways = 8! * 8P6 = 40320 * 20160 = 812851200